Can you explain in detail and clear please
Part A An infants toy has a 120 g wooden animal hanging from a spring. If pulled down gently, the animal oscillates up and down with a period of 0.51 s.His older sister pulls the spring a bit more than intended. She puls the animal 35 cm below its equilibrium position, then lets go. The animal fies upward and detaches from the spring right at the animals equilibrium position. the animal does not hit anything on the way up, how far above its equilibrium position will it go Express your answer with the appropriate units Value Units Submit
Answer #1

Angular speed,

w = 2*pi / T

w = 2*3.14 / 0.51 = 12.32 rad/s

Spring constant, k = mw^2

k = 0.12*(12.32)^2 = 18.21 N/m

From conservation of energy,

mgh = (1/2)kA^2

0.12*9.8*h = (1/2)*18.21*(0.35)^2

h = 0.948 m

h = 94.8 cm